eleminatio linear equation
3x + y=-1
2x +-y= 3
substitution
3x + 2y=1
7x + 6y=2

Answer by Alan3354(69443)

3x + y=-1
2x +-y= 3
--------------- Add
5x = 2
x = 2/5 etc.
--------------
substitution
3x + 2y=1 --> 9x + 6y = 3 --> 6y = 3 - 9x
7x + 6y=2
------------
Sub for 6y in Eqn 2
7x + (3-9x) = 2
-2x = -1
x = 1/2